OpenStudy (anonymous):

The graph of which function does not have a y-intercept of (0, 1)? A. y = -2^x B. y = 2^-x C. y =0.1^x D. y =0.5^x

OpenStudy (triciaal):

when x = 0 y = 1 put x = 0 in each equation which option does not give a value of 1?

OpenStudy (anonymous):

well these are the answer choices A. (0, 1) B. (1, 0) C. (1, 1) D. (0, 0)

OpenStudy (triciaal):

now I am confused. which graph does not have a y-intercept of (0, 1)? it would be graph A .when x = 0 then y is not equal to 1. why do you have these as the answer choices?

OpenStudy (anonymous):

My mistake the correct answer choices are A. y=-2^x B. y=2^-x C. y=0.1^x D. y=0.5^x

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

and read my response so you really understand why graph A

OpenStudy (anonymous):

Okay thanks @triciaal